OpenStudy (anonymous):

Solve: |m| = 8 Solve: |2t + 5| = 3 Solve: |x + 3| = 3 Solve: |3r - 12| = 0

OpenStudy (anonymous):

m=-8 or m=8 m=-4 or m=-1 x=-6 or x=0 r=4 , Isn't it?

OpenStudy (anonymous):

\[\Large \begin{array}{l} \left| m \right| = 8\\ m = \pm 8\\\\ \left| {2t + 5} \right| = 3\\ 2t + 5 = \pm 3\\ 2t = ( - 3 - 5),(3 - 5) = - 8, - 2\\ t = - 4, - 1\\\\ \left| {x + 3} \right| = 3\\ x + 3 = \pm 3\\ x = ( - 3 - 3),(3 - 3) = - 6,0\\\\ \left| {3r - 12} \right| = 0\\ 3r - 12 = 0\\ 3r = 12\\ r = 4 \end{array}\]

OpenStudy (anonymous):

m=8,-8 2t+5=3 ,2t+5=-3 2t=-2 , 2t=-8 t=-1 , t=-4 x+3=3 ,x+3=-3 x=0 ,x=-6 3r-12=0 3r=12 r=4